# Kerr enhanced optomechanical entanglement via reservoir design

### v1

@@ IV.2 The covariance matrix of the system
For the present linearized optomechanical system, its dynamics can be well described by both the first- and second-order moments. Considering that the first-order moments are initially zero, they will remain zero throughout the subsequent evolution, allowing us to neglect the first-order moments in the following analyses. In this section, we derive the equations of motion of the second-order moments and the steady-state covariance matrix of the system. In terms of the relation $\partial_{t}\langle\hat{o}_{i}\hat{o}_{j}\rangle=\text{Tr}(\dot{\hat{\rho}}_{sd}\hat{o}_{i}\hat{o}_{j})$ for $\hat{o}_{i},\hat{o}_{j}$in $\hat{a}$, $\hat{a}^{\dagger}$, $\hat{b}$, $\hat{b}^{\dagger}$} and based on the quantum master equation (III), we can obtain the equation of motion for these second-order moments of the system operators as
@@
where $\mathbf{u}_{i}$ for $i=$1-4 are the element of the operator vector $\mathbf{u}$ defined in Eq. (18). The elements of the covariance matrix can be further expressed as a function of the second-order moments [57]. In the steady state of the system, the covariance matrix takes the form as
@@
where the matrix $\mathbf{V_{O}}$ is associated with the optical mode, $\mathbf{V_{M}}$ is related to the mechanical resonator, and $\mathbf{C}$ characterizes their interaction, which quantifies the bipartite entanglement between the cavity field and the mechanical mode. To quantity the quantum entanglement for the Gaussian steady-state of the cavity field and mechanical mode, we employ the logarithm negativity, which is defined by [53, 54]
@@ IV.3 Enhanced optomechanical entanglement
We also investigate the influence of both the environmental thermal phonon occupation and optical driving amplitude on the entanglement generation performance. Figure 6(a) shows the logarithmic negativity versus the average thermal phonon occupation $\bar{n}_{th}$ and the Kerr constant $\chi$. It can be seen that, as $\chi$ increases, the entanglement initially increases and then decreases. For some values of $\chi$, even when the average thermal phonon occupation $\bar{n}_{th}$ reaches 3000, the steady-state optomechanical entanglement still remains. We point out that, though the $n_{th}$ is very large, the average phonon number in the mechanical mode is small. This is because the working point of the system is right useful for the sideband cooling, and hence the ground-state cooling of the mechanical resonator can be realized. We checked the average phonon number in the mechanical mode and find that $\langle b^{\dagger}b\rangle_{ss}\approx 0.5$. We also find that, for a given $n_{th}$, the $E_{N}$ first increases and then decreases with the increase of $\chi$. This implies that the introduction of $\chi$ improves the ability of the system to resist thermal noise. This phenomenon can be explained based on the dependence of $G_{sd}$ on the Kerr constant $\chi$. Compared to the case of $\chi=0$, the coupling strength $G_{sd}/\omega_{m}$ is largely enhanced for a wide range of $\Delta_{c}/\omega_{m}$. In Fig. 6(b), we show the logarithmic negativity versus the optical driving amplitude $\Omega$ and Kerr constant $\chi$. It can be observed that as the driving amplitude increases, the entanglement first increases and then decreases, suggesting that stronger driving does not always lead to greater entanglement. Similarly, the peak value of $E_{N}$ appears around $\chi/\omega_{m}\approx 2.4\times 10^{-5}$.
@@ V Discussions and Conclusion
Finally, we present some discussions on the experimental implementation of this scheme. In our scheme, there are four elements: the optomechanical coupling between the field and the mechanical motion, the Kerr interaction of the cavity field, the designed squeezed vacuum bath of the cavity field, and the monochromatic driving of the cavity field. To implement the scheme, the physical system candidate should be able to realize these four elements. The optomechanical coupling can now be implemented in many physical systems, including optical microresonators [58, 59, 60], electromechanical systems [61, 62, 63, 64], photonic crystal nanobeams [65, 66, 67], and Fabry-Pérot cavities [68]. The monochromatic driving for the cavity field is experimentally accessible in many physical systems. Therefore, the key elements are the Kerr interaction and the squeezed vacuum bath. Usually, the Kerr interaction can be induced by coupling the cavity field to an $N$-type four-level atomic system. It has been reported that the Kerr constant can reach $\chi/2\pi\approx 0.16$ MHz [48]. In addition, we point out that the Kerr interaction of the cavity field exists in the circuit-QED system [69] and the Kerr interaction has been realized in cavity magnomechanical systems [70, 71]. In particular, several experiments have reported the realization of the Kerr interaction [72, 73, 74]. For the squeezed vacuum bath, it can be created by injecting a squeezed vacuum field into the cavity mode. Based on the above analyses, our scheme should be within the reach of current and near-future experimental conditions.
In conclusion, we have proposed a scheme to enhance the optomechanical entanglement. This is achieved by introducing both the Kerr interaction in the cavity field and the squeezed vacuum bath. Under the strong driving of the cavity field, the two-photon terms induced by the Kerr interaction will introduce an effective amplification into the linearized optomechanical coupling. Meanwhile, the optical bath will also be amplitied by the two-photon terms and this amplification effect can be eliminated by introducing a well designed squeezed vacuum reservoir. We have analyzed the parameter space and discussed the stability of the linearized optomechanical system. We have also studied the enhanced optomechanical entanglement. It has been found that the optomechanical entanglement can be enhanced with the coupling amplification induced by the Kerr nonlinearity. Our results will motivate the study on enhancing other optomechanical effects via quantum amplification and reservoir design.
@@ Acknowledgements.
The authors thank for Dr. Jian Huang and Dr. Deng-Gao Lai for helpful discussions. J.-Q.L. was supported in part by National Natural Science Foundation of China (Grants No. 12175061, No. 12247105, No. 11935006, and No. 12421005), National Key Research and Development Program of China (Grant No. 2024YFE0102400), and Hunan Provincial Major Sci-Tech Program (Grant No. 2023ZJ1010).
@@ Appendix A Analyses of the stability of the system
[…]
where $z=\exp(2i\pi/3)$, and we introduce $\eta=q^{2}/3+p^{3}/3$ with $p=(3ac-b^{2})/(3a^{2})$ and $q=(-2b^{3}+9abc-27a^{2}d)/(27a^{3})$.
@@
[…]
@@
where “$\det(\cdot)$” denotes the determinant of a matrix. Then the Routh matrix can be expressed as

### v2

@@ IV.2 The covariance matrix of the system
For the present linearized optomechanical system, its dynamics can be well described by both the first- and second-order moments. Considering that the first-order moments are initially zero, they will remain zero throughout the subsequent evolution, allowing us to neglect the first-order moments in the following analyses. In this section, we derive the equations of motion of the second-order moments and the steady-state covariance matrix of the system. In terms of the relation $\partial_{t}\langle\hat{o}_{i}\hat{o}_{j}\rangle=\text{Tr}(\dot{\hat{\rho}}_{sd}\hat{o}_{i}\hat{o}_{j})$ for $\hat{o}_{i},\hat{o}_{j}$ $\in$ $\hat{a}$, $\hat{a}^{\dagger}$, $\hat{b}$, $\hat{b}^{\dagger}$} and based on the quantum master equation (III), we can obtain the equation of motion for these second-order moments of the system operators as
@@
where $\mathbf{u}_{i}$ for i=i=1-4 are the element of the operator vector $\mathbf{u}$ defined in Eq. (18). The elements of the covariance matrix can be further expressed as a function of the second-order moments [69]. In the steady state of the system, the covariance matrix takes the form
@@
where the matrix $\mathbf{V_{O}}$ is associated with the optical mode, $\mathbf{V_{M}}$ is related to the mechanical resonator, and $\mathbf{C}$ characterizes their interaction, which quantifies the bipartite entanglement between the cavity field and the mechanical mode. To quantity the quantum entanglement for the Gaussian steady state of the cavity field and mechanical mode, we employ the logarithm negativity, which is defined by [64, 65]
@@ IV.3 Enhanced optomechanical entanglement
We also investigate the influence of both the environmental thermal phonon occupation and optical driving amplitude on the entanglement generation performance. Figure 6(a) shows the logarithmic negativity versus the average thermal phonon occupation $\bar{n}_{th}$ and the Kerr constant $\chi$. It can be seen that, as $\chi$ increases, the entanglement initially increases and then decreases. For some values of $\chi$, even when the average thermal phonon occupation $\bar{n}_{th}$ reaches 3000, the steady-state optomechanical entanglement still remains. We point out that athough the $n_{th}$ is very large, the average phonon number in the mechanical mode is small. This is because the working point of the system is very useful for the sideband cooling, and hence the ground-state cooling of the mechanical resonator can be realized. We checked the average phonon number in the mechanical mode and find that $\langle b^{\dagger}b\rangle_{ss}\approx 0.5$. We also find that, for a given $n_{th}$, the $E_{N}$ first increases and then decreases with the increase of $\chi$. This implies that the introduction of $\chi$ improves the ability of the system to resist thermal noise. This phenomenon can be explained based on the dependence of $G_{sd}$ on the Kerr constant $\chi$. Compared to the case of $\chi=0$, the coupling strength $G_{sd}/\omega_{m}$ is largely enhanced for a wide range of $\Delta_{c}/\omega_{m}$. In Fig. 6(b), we show the logarithmic negativity versus the optical driving amplitude $\Omega$ and Kerr constant $\chi$. It can be observed that as the driving amplitude increases, the entanglement first increases and then decreases, suggesting that stronger driving does not always lead to greater entanglement. Similarly, the peak value of $E_{N}$ appears around $\chi/\omega_{m}\approx 2.4\times 10^{-5}$.
Figure 7: Logarithmic negativity $E_{\mathcal{N}}$ as a function of $\chi/\omega_{m}$ for different values of the parameters $(r_{e},\theta_{e})$, under the four cavity damping rates: (a) $\kappa_{a}/\omega_{m}=0.5$, (b) $\kappa_{a}/\omega_{m}=0.8$, (c) $\kappa_{a}/\omega_{m}=1.2$, and (d) $\kappa_{a}/\omega_{m}=1.5$. Each panel includes nine curves for $r_{e}=\{0,r/2,r\}$ and $\theta_{e}=\{0+\pi,\varphi/2+\pi,\varphi+\pi\}$, illustrating the impact of the squeezed vacuum reservoir on the steady-state optomechanical entanglement. Here, $\Delta_{c}/\omega_{m}=0$, and other parameters are the same as Fig. 2.
In the above analyses, we demonstrate the enhanced optomechanical entanglement based on Eq. (16), where the two-photon noise terms induced by the Kerr interaction are completely canceled by the introduced squeezed reservoir [the parameters $N_{ss}$ and $M_{ss}$ in Eqs. (14) are equal to zero]. Next, we investigate the optomechanical entanglement for the case where the amplitude or phase of the squeezed reservoir does not give rise to perfect cancellation of these two-photon noise terms. To this end, we numerically calculate the logarithm negativity to characterize the optomechanical entanglement for various amplitudes $r_{e}$ and phases $\theta_{e}$ under these four cavity damping rates $\kappa_{a}/\omega_{m}=0.5,0.8,1.2$, and $1.5$, as shown in Fig. 7.
In Fig. 7, we can see that the logarithm negativity reaches its maximum for the parameter $(r_{e},\theta_{e})=(r,\varphi+\pi)$, while it is very small for the parameter $(r_{e},\theta_{e})=(0,0+\pi)$. The results mean that the optomechanical entanglement is optimized when the two-photon term induced by the Kerr interaction is completely canceled by the squeezed vacuum reservoir, while the absence of such a reservoir leads to a strong suppression for the optomechanical entanglement. Meanwhile, the intermediate configurations such as $(r_{e},\theta_{e})=(0,\varphi/2+\pi),(0,\varphi+\pi),(r/2,0+\pi),(r/2,\varphi/2+\pi),(r/2,\varphi+\pi),(r,0+\pi),$ and $(r,\varphi/2+\pi)$ clearly show reduced logarithm negativity, which indicate that, though the optomechanical entanglement is sensitive to mismatches in the squeezing parameters $r_{e}$ and $\theta_{e}$, the system still tolerates moderate deviations. This highlights the experimental feasibility of the proposed scheme, even with the imperfect squeezed vacuum reservoir.
@@ V Discussions and Conclusion
Finally, we present some discussions on the experimental implementation of this scheme. In our scheme, there are four elements: the optomechanical coupling between the field and the mechanical motion, the Kerr interaction of the cavity field, the designed squeezed vacuum bath of the cavity field, and the monochromatic driving of the cavity field. To implement the scheme, the physical system candidate should be able to realize these four elements. The optomechanical coupling can now be implemented in many physical systems, including optical microresonators [70, 71, 72, 73], electromechanical systems [74, 75, 76, 77, 78], photonic crystal nanobeams [79, 80], and Fabry-Pérot cavities [81]. The monochromatic driving for the cavity field is experimentally accessible in many physical systems. Therefore, the key elements are the Kerr interaction and the squeezed vacuum bath. Usually, the Kerr interaction can be induced by coupling the cavity field to an $N$-type four-level atomic system. It has been reported that the Kerr constant can reach $\chi/2\pi\approx 0.16$ MHz [59]. In addition, we point out that the Kerr interaction of the cavity field exists in the circuit-QED system [82] and the Kerr interaction has been realized in cavity magnomechanical systems [83, 84]. In particular, several experiments have reported the realization of the Kerr interaction [85, 86, 87]. Notably, recent experiments have demonstrated Kerr-enhanced optomechanical interactions in both microwave and optical systems. In particular, superconducting circuit electromechanics [88] and optical Fabry-Pérot cavities at 1064 nm [58] both show the compatibility of the Kerr effect with optomechanical platforms. For the squeezed vacuum bath, it can be created by injecting a squeezed vacuum field into the cavity mode. Based on the above analyses, our scheme should be within the reach of current and near-future experimental conditions.
In conclusion, we have proposed a scheme to enhance the optomechanical entanglement. This is achieved by introducing both the Kerr interaction in the cavity field and the squeezed vacuum bath. Under the strong driving of the cavity field, the two-photon terms induced by the Kerr interaction will introduce an effective amplification into the linearized optomechanical coupling. Meanwhile, the optical bath will also be amplitied by the two-photon terms and this amplification effect can be eliminated by introducing a well designed squeezed vacuum reservoir. We have analyzed the parameter space and discussed the stability of the linearized optomechanical system. We have also studied the enhanced optomechanical entanglement. It has been found that the optomechanical entanglement can be enhanced with the coupling amplification induced by the Kerr nonlinearity. Our results will motivate the study of enhancing other optomechanical effects via quantum amplification and reservoir design.
@@ Acknowledgements.
The authors thank Dr. Jian Huang and Dr. Deng-Gao Lai for helpful discussions. J.-Q.L. was supported in part by the National Natural Science Foundation of China (Grants No. 12175061, No. 12575015, No. 12247105, No. 11935006, and No. 12421005), National Key Research and Development Program of China (Grant No. 2024YFE0102400), and Hunan Provincial Major Sci-Tech Program (Grant No. 2023ZJ1010). C.L. was supported in part by the Postgraduate Scientific Research Innovation Project of Hunan Province (Grant No. CX20250721) and the Tenglong Innovative Talent Fund of Hunan Normal University (Grant No. 2025TL209).
@@ Appendix A Analyses of the stability of the system
$y_{n=1,2,3}=-\frac{b}{3a}+z^{n}\sqrt[3]{\frac{q}{2}+\sqrt{\eta}}+z^{2n}\sqrt[3]{\frac{q}{2}-\sqrt{\eta}},$~ (38)
where $z=\exp(2i\pi/3)$, and we introduce $\eta=(q/2)^{2}+(p/3)^{3}$ with $p=(3ac-b^{2})/(3a^{2})$ and $q=(-2b^{3}+9abc-27a^{2}d)/(27a^{3})$.
@@
$y_{3}=-\frac{b}{3a}+\sqrt[3]{\frac{q}{2}+\sqrt{\eta}}+\sqrt[3]{\frac{q}{2}-\sqrt{\eta}}.$~ (39)
@@
where $\det(\cdot)$” denotes the determinant of a matrix. Then the Routh matrix can be expressed as